# Seeking Consistent Flat Minima for Better Domain Generalization via Refining Loss Landscapes

### v1

@@ Abstract
Domain generalization aims to learn a model from multiple training domains and generalize it to unseen test domains. Recent theory has shown that seeking the deep models, whose parameters lie in the flat minima of the loss landscape, can significantly reduce the out-of-domain generalization error. However, existing methods often neglect the consistency of loss landscapes in different domains, resulting in models that are not simultaneously in the optimal flat minima in all domains, which limits their generalization ability. To address this issue, this paper proposes an iterative Self-Feedback Training (SFT) framework to seek consistent flat minima that are shared across different domains by progressively refining loss landscapes during training. It alternatively generates a feedback signal by measuring the inconsistency of loss landscapes in different domains and refines these loss landscapes for greater consistency using this feedback signal. Benefiting from the consistency of the flat minima within these refined loss landscapes, our SFT helps achieve better out-of-domain generalization. Extensive experiments on DomainBed demonstrate superior performances of SFT when compared to state-of-the-art sharpness-aware methods and other prevalent DG baselines. On average across five DG benchmarks, SFT surpasses the sharpness-aware minimization by 2.6% with ResNet-50 and 1.5% with ViT-B/16, respectively. The code will be available soon.
@@ 2 Related Work
Loss landscapes for model generalization. The idea of searching for “flat” minima of loss landscapes can be traced back to Hochreiter and Schmidhuber [26], who proposed penalizing the sharpness of minima as a form of regularization, related to the Minimum Description Length (MDL) principle. Over the past few decades, numerous studies [27, 28, 29] have linked the flatness of loss landscapes to the generalization ability of deep models. As a result, many optimization methods have been developed to find flatter minima, such as Entropy-SGD [32] and the diffusion approach [31]. More recently, two methods, Stochastic Weight Averaging (SWA) [71] and Sharpness-Aware Minimization (SAM) [34], have gained significant attention for their effectiveness and scalability. SWA finds flatter minima by averaging model parameters across the optimization trajectory, while SAM seeks parameters within regions of the loss landscape that exhibit uniformly low loss values. Jean Kaddour et al.  [72] provided a detailed comparison of these methods, showing that SAM tends to lead models to wider basins, while SWA merely helps locate the center of one certain basin. In this paper, we focus on addressing the inconsistency of loss landscapes when SAM is applied in domain generalization (DG) scenarios.
Loss landscapes for domain generalization. Recently, optimization techniques aiming at finding flat minima in loss landscapes have also been applied to improve the out-of-domain generalization of deep models [30]. Junbum Cha et al. [70] introduced the flatness to the domain generalization and proposed a novel Stochastic Weight Averaging Densely (SWAD), which is an extension of SWA with a dense and overfit-aware stochastic weight sampling strategy. Zhang et al. [103] pointed out that SAM-based methods are limited to zeroth-order flatness, which may be insufficient to discriminate minima with low generalization error. Thus, they developed a novel method named GAM, which targets first-order flatness to enhance generalization. Moreover, FAD [104] optimizes both zeroth- and first-order flatness simultaneously for domain generalization. However, these methods primarily focus on seeking better flat minima for DG, ignoring the influence of landscape discrepancy across multiple domains. In contrast, Giambattista Parascandolo et al. [56] considered the domain discrepancy via formalizing a notion of consistency for minima of loss surfaces, and then proposed an algorithm (ANDMask) to alter the optimization trajectory through masking inconsistent gradient components among training domains. Different from all of the above methods that consider loss landscapes to be static, we attempt to refine the loss surfaces dynamically for consistency and finally seek a consistent flat minima across various domains.
@@ 6      The Feedback Phase:
8      Update the model parameter $\boldsymbol{\theta}$ on the domain $D_{d}$:
9      $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}_{D_{d}}^{\mathrm{SAM-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})$.
@@ 3.2.2 The Refinement Phase
To ensure label correctness, we could use the standard cross-entropy loss to force the generated soft labels to approximate one-hot labels. However, this approach carries the risk of overfitting to the one-hot labels, leading to overconfidence and limiting the model’s generalization ability. One technique that attempts to mitigate this issue is label smoothing [39], but it merely introduces noise to the one-hot labels, making them less informative and unable to capture the underlying relationships between categories and different input samples. This paper focuses on a more fundamental problem: minimizing the “distance” between the refiner’s output and the true label space for each category. For instance, if the first category is the true label for the sample $\boldsymbol{x}_{i}$, the corresponding label space is defined as:
@@
$\min_{\mathbf{y}}\,\mathrm{KL}(\mathbf{y}\,||\,\tilde{\mathbf{y}})\quad\text{subject to}\quad\mathbf{y}\in C_{1}.$ (11)
We propose an efficient algorithm (Algorithm 2) to find the optimal solution $\mathbf{y}^{\star}$ for this problem, which converges in $N$ steps ($N$ is the number of classes) and is computationally more efficient than using common convex programming tools. For the condition that $y_{i}\neq 1$, we can also get the optimal solution via category swapping before and after applying Algorithm 2. Once having $\mathbf{y}^{\star}$, the task becomes minimizing the KL divergence between $\mathbf{y}^{\star}$ and $\tilde{\mathbf{y}}$: $\mathrm{KL}(\mathbf{y}^{\star}\,||\,\tilde{\mathbf{y}})$. Since $\mathbf{y}^{\star}$ depends on $\boldsymbol{\phi}$ and Algorithm 2 may involve non-differentiable operations, we adopt a widely-used alternative iterative strategy: we treat $\mathbf{y}^{\star}$ as constant during each update step for $\boldsymbol{\phi}$, and after updating $\boldsymbol{\phi}$, we recalculate $\mathbf{y}^{\star}$ using Algorithm 2. Minimizing the KL divergence then leads to an objective similar to cross entropy:
@@ 13             Break
To encourage the loss landscapes to become flatter while pursuing consistency, we add the the sharpness term to the previously obtained feedback signal (the sharpness difference) as an extra penalty. Then, the final loss function is defined as:
@@ 4.2.2 Main Results
Comparison with weight average methods. Weight averaging methods, such as SWAD [70], seek flat minima in a different manner. Previous work [72] has compared it with SAM and revealed that parameters sought by the two methods are located in different basins. Since model averaging is a totally different method of model selection, it wouldn’t be fair to make a direct comparison with SFT. Nonetheless, in our experiment with ViT-B/16 (as shown in Table 4), we observed that SAM and SFT outperform SWAD by 0.6% (85.1$\rightarrow$85.7%) and 1.4% (85.1$\rightarrow$86.5%), respectively.
@@ 4.3 Ablation Study
Table 4: Comparisons of SFT and the weight average method (SWAD). We report out-of-domain accuracies with/without SWAD on four domains of OfficeHome.
Algorithms Art Clipart Product Real World Avg.
ERM [45] 83.7 73.8 89.9 89.2 84.1
ERM [45]$+$SWAD [70] 84.9 74.9 90.8 89.9 85.1
SAM [34] 84.2 76.6 91.0 90.8 85.7
SAM [34]$+$SWAD [70] 84.6 76.5 91.2 90.8 85.8
SFT (Ours) 86.3 77.7 91.1 90.9 86.5
SFT (Ours) $+$ SWAD [70] 86.0 77.5 91.2 91.3 86.5
@@ 10.1 Code
Our work is built upon DomainBed, which is released under the MIT license. All experiments are conducted on a single NVIDIA Tesla V100 or A40. Later, we will upload our code, along with the log files containing additional experimental details, to GitHub.
@@ 10.2.2 Experiments on Real Dataset
The hyperparameter search spaces for ResNet-50, ViT-B/16 and ViT-L/14 are shown below. In the table, $U$ and list indicate Uniform distribution and random choice, respectively. The optimal hyperparameters and model parameters for each dataset will be available on our GitHub page.

### v2

@@ Abstract
Domain generalization aims to learn a model from multiple training domains and generalize it to unseen test domains. Recent theory has shown that seeking the deep models, whose parameters lie in the flat minima of the loss landscape, can significantly reduce the out-of-domain generalization error. However, existing methods often neglect the consistency of loss landscapes in different domains, resulting in models that are not simultaneously in the optimal flat minima in all domains, which limits their generalization ability. To address this issue, this paper proposes an iterative Self-Feedback Training (SFT) framework to seek consistent flat minima that are shared across different domains by progressively refining loss landscapes during training. It alternatively generates a feedback signal by measuring the inconsistency of loss landscapes in different domains and refines these loss landscapes for greater consistency using this feedback signal. Benefiting from the consistency of the flat minima within these refined loss landscapes, our SFT helps achieve better out-of-domain generalization. Extensive experiments on DomainBed demonstrate superior performances of SFT when compared to state-of-the-art sharpness-aware methods and other prevalent DG baselines. On average across five DG benchmarks, SFT surpasses the sharpness-aware minimization by 2.6% with ResNet-50 and 1.5% with ViT-B/16, respectively.
@@ 2 Related Work
Loss landscapes for model generalization. The idea of searching for “flat” minima of loss landscapes can be traced back to Hochreiter and Schmidhuber [21], who proposed penalizing the sharpness of minima as a form of regularization, related to the Minimum Description Length (MDL) principle. Over the past few decades, numerous studies [30, 14, 27] have linked the flatness of loss landscapes to the generalization ability of deep models. As a result, many optimization methods have been developed to find flatter minima, such as Entropy-SGD [11] and the diffusion approach [45]. More recently, two methods, Stochastic Weight Averaging (SWA) [23] and Sharpness-Aware Minimization (SAM) [16], have gained significant attention for their effectiveness and scalability. SWA finds flatter minima by averaging model parameters across the optimization trajectory, while SAM seeks parameters within regions of the loss landscape that exhibit uniformly low loss values. Jean Kaddour et al.  [28] provided a detailed comparison of these methods, showing that SAM tends to lead models to wider basins, while SWA merely helps locate the center of one certain basin. Several efficient variants of SAM [12, 26, 24] have also been developed for real-world applications. In this paper, we focus on addressing the inconsistency of loss landscapes when SAM is applied in DG scenarios.
Loss landscapes for domain generalization. Recently, optimization techniques aiming at finding flat minima in loss landscapes have also been applied to improve the out-of-domain generalization of deep models [73]. Junbum Cha et al. [9] introduced the flatness to the domain generalization and proposed a novel Stochastic Weight Averaging Densely (SWAD), which is an extension of SWA with a dense and overfit-aware stochastic weight sampling strategy. Zhang et al. [68] pointed out that SAM-based methods are limited to zeroth-order flatness, which may be insufficient to discriminate minima with low generalization error. Thus, they developed a novel method named GAM, which targets first-order flatness to enhance generalization. Moreover, FAD [67] optimizes both zeroth- and first-order flatness simultaneously for domain generalization. However, these methods primarily focus on seeking better flat minima for DG, ignoring the influence of landscape discrepancy across multiple domains. In contrast, Giambattista Parascandolo et al. [48] considered the domain discrepancy via formalizing a notion of consistency for minima of loss surfaces, and then proposed an algorithm (ANDMask) to alter the optimization trajectory through masking inconsistent gradient components among training domains. Different from all of the above methods that consider loss landscapes to be static, we attempt to refine the loss surfaces dynamically for consistency and finally seek consistent flat minima across various domains.
@@ 6      The Feedback Phase:
8      Update the model parameter $\boldsymbol{\theta}$ on the domain $D_{d}$: $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}_{D_{d}}^{\mathrm{SAM-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})$.
@@ 3.2.2 The Refinement Phase
To ensure label correctness, we could use the standard cross-entropy loss to force the generated soft labels to approximate one-hot labels. However, this approach carries the risk of overfitting to the one-hot labels, leading to overconfidence and limiting the model’s generalization ability. One technique that attempts to mitigate this issue is label smoothing [46], but it merely introduces noise to the one-hot labels, making them less informative and unable to capture the underlying relationships between categories and different input samples. This paper focuses on a more fundamental problem: minimizing the “distance” between the refiner’s output and the true label space for each category. For instance, if the first category is the true label for the sample $\boldsymbol{x}_{i}$ (i.e., $y_{i}=1$), the corresponding label space is defined as:
$C_{1}=\{(q_{1},\ldots,q_{N})\mid\forall k\neq 1:q_{1}\geq\alpha q_{k},\sum_{k=1}^{N}q_{k}=1\},$ (10)
@@
where $\mathrm{KL}(\cdot\|\cdot)$ denotes the Kullback-Leibler (KL) divergence. We propose an efficient algorithm (Algorithm 2) to find the optimal solution $\mathbf{y}^{\star}$ for this problem, which converges in $N$ steps ($N$ is the number of classes) and is computationally more efficient than using common convex programming tools. For the condition that $y_{i}\neq 1$, we can also get the optimal solution via category swapping before and after applying Algorithm 2. Once having $\mathbf{y}^{\star}$, the task becomes minimizing the KL divergence between $\mathbf{y}^{\star}$ and $\tilde{\mathbf{y}}$: $\mathrm{KL}(\mathbf{y}^{\star}\,||\,\tilde{\mathbf{y}})$. Since $\mathbf{y}^{\star}$ depends on $\boldsymbol{\phi}$ and Algorithm 2 may involve non-differentiable operations, we adopt a widely-used alternative iterative strategy: we treat $\mathbf{y}^{\star}$ as constant during each update step for $\boldsymbol{\phi}$, and after updating $\boldsymbol{\phi}$, we recalculate $\mathbf{y}^{\star}$ using Algorithm 2. Minimizing the KL divergence then leads to an objective similar to cross entropy:
@@ 13             Break
To encourage the loss landscapes to become flatter while pursuing consistency, we add the sharpness term to the previously obtained feedback signal (the sharpness difference) as an extra penalty. Then, the final loss function is defined as:
@@ 4.2.2 Main Results
4.3 Sharpness Analysis
We also perform a quantitative analysis of the landscape consistency. As mentioned earlier, SFT quantifies landscape consistency through domain-specific sharpness differences. In Figure 4, we find that SFT reduces domain-specific sharpness values and their differences compared to SAM, while simultaneously attaining higher classification accuracy. These empirical results confirm that SFT effectively improves the loss landscape consistency across domains, while also demonstrating the positive impact of landscape consistency on the DG performance.
@@ 4.4 Ablation Study
Figure 4: Comparison of model sharpness and DG accuracy across different training strategies on the OfficeHome dataset.
@@ 10.1 Code
Our work is built upon DomainBed, which is released under the MIT license. All experiments are conducted on a single NVIDIA Tesla V100 or A40.
@@ 10.2.2 Experiments on Real Dataset
The hyperparameter search spaces for ResNet-50, ViT-B/16 and ViT-L/14 are shown below. In the table, $U$ and list indicate Uniform distribution and random choice, respectively.